A long straight wire carries a current. The strength of the magnetic field, due to the wire, 10.0 cm from the axis of the wire is 20T . Determine the current in the wire. 


1
Expert's answer
2022-01-16T13:09:43-0500

"I=\frac{2\pi Br}{\mu_0}=\frac{2\pi\cdot 20\cdot10^{-6}\cdot0.1}{4\pi\cdot 10^{-7}}=10~A."